PRESSURE BROADENING AND FREQUENCY MEASUREMENTS OF NITRIC ACID LINES IN THE 683 GHZ REGION

\underline{L. ZU}, P. B. DAVIES, Department of Chemistry, University of Cambridge, Cambridge CB2 1EW, UK; P. A. HAMILTON, Queen Mary and Westfield College, London E1 4NS, UK.

Nitrogen, oxygen and self-broadening coefficients of four transitions in the ground vibrational state of nitric acid (HNO_{3}) have been measured at room temperature, using tunable far infrared (TuFIR) laser spectroscopy. In addition, absolute frequencies of 60 HNO_{3} lines between 676 and 690 GHz were measured for the first time with a mean uncertainty of < 1 MHz. These results provide a significant database for atmospheric remote sensing experiments and spectral analyses in this region.